Assume that the

demand for good X has a for that can be expressed by De following equation: X = 5000 - 20p- + 100py - 2L According is a research we know that

demanded the price of

good X and L is 10 and the income is 400. can be said that in lems

of income elashoy the good is a normal good?


1
Expert's answer
2021-10-26T17:22:14-0400

Solution:

First, calculate the income elasticity of demand:

Find the value of X:

X = 5000 – 20(10) + 100(400) – 2(10) = 5000 – 200 + 40,000 – 20 = 44,780

X = 44,780

Income elasticity of demand (YED) = "\frac{\triangle X}{\triangle I} \times \frac{I}{X}"

"\frac{\triangle X}{\triangle I} = 100"


YED = 100 * 400/44,780 = 0.89


Income elasticity of demand (YED) = 0.89


YED > 0 and therefore a normal good.

A normal good has a demand elasticity of income greater than zero. This means that the demand for a normal good will rise as consumer income rises.